# Emergent Hawking Radiation and Quantum Sensing in a Quenched Chiral Spin Chain

### v1

@@ Abstract
We investigate the emergence and detection of Hawking radiation (HR) in a 1D chiral spin chain model, where the gravitational collapse is simulated by a sudden quantum quench that triggers a horizon-inducing phase transition. While our previous work Jaiswal and Shankaranarayanan [2025] established that this model mimics BH formation conditions even when the Hoop conjecture is seemingly violated, we here focus on the resulting stationary radiation spectrum and its detectability. By mapping the spin chain dynamics to a Dirac fermion in a curved $(1+1)$-dimensional spacetime, we analyze the radiation using two complementary approaches: field-theoretic modes and operational quantum sensors. First, using localized Gaussian wave packets to model realistic detectors, we find that the radiation spectrum exhibits deviations from the ideal Planckian form, analogous to frequency-dependent greybody factors, while retaining robust Poissonian statistics that signal the loss of formation-scale information. Second, we introduce a qubit coupled to the chain as a stationary Unruh-DeWitt detector. We demonstrate that the qubit functions as a faithful quantum sensor of the Hawking temperature only in the weak-coupling regime, where its population dynamics are governed solely by the bath spectral density. In the strong-coupling limit, the probe thermalizes with the global environment, obscuring the horizon-induced thermal signature. These results provide a clear operational protocol for distinguishing genuine analog HR from environmental noise in quantum simulation platforms.
@@
Introduction: A central challenge in semiclassical and quantum gravity is the direct experimental verification of Hawking radiation (HR) Parker [1975], Corley and Jacobson [1996], Shankaranarayanan et al. [2002, 2001], Casadio [2002], Kumar and Shankaranarayanan [2016], Pal et al. [2022], Smerlak and Singh [2013], Pal and Fischer [2024]. While astrophysical observations, such as strong-field gravitational lensing, provide critical probes of the near-horizon geometry of black holes (BHs) Mandal and Shankaranarayanan [2026], Virbhadra [2022], the quantum dynamics of the horizon itself remain obscured by the weakness and delocalization of the radiation. Analogue gravity platforms have emerged as a powerful tool, as they enable controlled access to near-horizon quantum effects in laboratory settings ranging from fluids to optical systems Barcelo et al. [2005], Steinhauer [2014], Steinhauer and de Nova [2017], Banerjee et al. [2019], Guo [2022].
@@
To extract these elusive signatures from the background of a many-body system, we turn to the paradigm of quantum sensing Degen et al. [2017], Pirandola et al. [2018], DeMille et al. [2024], Mercer et al. [2025], Zhang and Zhuang [2021]. Controlled spin-1/2 particles (qubits) have established themselves as highly sensitive probes of complex quantum environments, underpinning modern sensing technologies across molecular, atomic, and solid-state systems Degen et al. [2017]. When coupled to an external bath, a qubit’s decoherence dynamics and population relaxation are governed by the spectral properties of the environment Płodzień et al. [2025]. This allows for the precise extraction of noise spectra Clerk et al. [2010], Young and Whaley [2012] and Hamiltonian parameters Degen et al. [2017] solely through measurements on the two-level probe.
@@
Motivated by these capabilities, we employ a qubit as a stationary detector — analogous to an Unruh-DeWitt detector — to probe the analogous radiation in the chiral spin chain. In our previous work, we established that the chiral spin chain Hamiltonian can mimic gravitational collapse, satisfying black hole formation conditions even when the Hoop conjecture is violated Jaiswal and Shankaranarayanan [2025]. In the present work, we focus on the implications for HR. Specifically, we investigate how the qubit’s decoherence profile encodes the emergent thermal features of the horizon.
@@
Our analysis proceeds in two steps. First, we examine the radiation spectrum directly using both idealized plane waves and physically realistic Gaussian wave packets. We show that while plane waves recover the standard thermal spectrum, localized packets reveal deviations akin to greybody factors Unruh [1976a], Page [1976], challenging the notion of HR as information-free. Second, we determine the operational conditions for detecting this temperature. We demonstrate that a qubit acts as a faithful thermometer only in the weak-coupling regime, whereas strong coupling leads to thermalization with the bulk environment rather than the horizon. Finally, we discuss the statistical nature of the emission, showing that the radiation remains Poissonian regardless of the geometric “Hoop” history, suggesting a universal erasure of formation scale information in the Hawking regime.
@@
Model setup: To explore the dynamics of HR, we model the gravitational collapse as a sudden quantum quench Byju et al. [2023], Kaplanek and Burgess [2021], Bongs et al. [2023], Covey et al. [2025], Balatsky et al. [2025]. We model the transition from a pre-horizon mass distribution to a BH using a time-dependent Hamiltonian $H(t)$. We assume the system initially evolves under a background Hamiltonian $H_{0}$ (representing the pre-collapse phase) and undergoes a sudden change at $t=t_{0}$ representing the formation of the horizon:
@@
where $\Theta(t-t_{0})$ is the Heaviside step function and $H_{h}$ denotes horizon-inducing dynamics and the probe interaction. In our analogue model, the pre-horizon regime is governed by the standard $XX$ spin chain, given by Jaiswal and Shankaranarayanan [2025]:
$H_{0}\equiv H_{XX}=-\sum_{j=1}^{N}(\mathcal{U}/{2})\left(\sigma_{j}^{x}\sigma_{j+1}^{x}+\sigma_{j}^{y}\sigma_{j+1}^{y}\right),$ (2)
where $\sigma_{j}^{\alpha}$ are Pauli operators at site $j$ satisfying periodic boundary conditions, and $\mathcal{U}$ is the nearest-neighbor coupling strength. For $t\geq t_{0}$, the BH formation is captured by the activation of the chirality term $H_{\chi}$ and the coupling to an external probe $H_{I}$. The quench Hamiltonian is defined as $H_{h}=H_{\chi}+H_{I}$. The chirality term is:
$H_{\chi}=\sum_{j=1}^{N}(v/4)\,\chi_{j},$ (3)
where $\chi_{j}=\vec{\sigma}_{j}\cdot(\vec{\sigma}_{j+1}\times\vec{\sigma}_{j+2})$ is the three-spin chirality operator Horner et al. [2023], Forbes et al. [2023] and $v$ is a dimensionless coupling constant. To detect the thermal signatures associated with HR, we couple the chain to a qubit subsystem (a two-level system) via the interaction term Quan et al. [2006], Liu et al. [2009], Jaiswal et al. [2022]:
$H_{I}=\sum_{j=1}^{N}(g_{1}\sigma_{q}^{z}+g_{2}\sigma_{q}^{x})\chi_{j}/4,$ (4)
where $\sigma_{q}^{x,z}$ act on the qubit and $g_{1,2}$ are coupling strengths. As shown in Fig. 1, coupling allows the qubit to exchange energy with the chain, inducing decoherence that serves as a proxy for the thermal bath of HR.
Figure 1: Schematic diagram of the Qubit-Environment Coupling in Chiral Spin Chain.
@@
In each sector corresponding to eigenvalues $\pm\sqrt{g_{1}^{2}+g_{2}^{2}}$, the spin chain evolves under an effective Hamiltonian. Following the standard Jordan-Wigner and Fourier transformations Jaiswal and Shankaranarayanan [2025], Jaiswal et al. [2021], the total Hamiltonian is diagonalized in momentum space (where $p\in[-\pi,\pi)$):
@@
$E_{\pm}(p)=-2\mathcal{U}\cos(p)+V_{e}^{\pm}\sin(2p).$ (7)
Here, the effective chiral coupling is modified by the probe interaction: $V_{e}^{\pm}=v\pm\sqrt{g_{1}^{2}+g_{2}^{2}}$.
Quantum Phase Transition and Time Evolution: Analysis of the ground-state energy density reveals a Quantum Phase Transition (QPT) driven by the competition between the nearest-neighbor hopping and the chiral interaction. A discontinuity in the second derivative of the energy density occurs at the critical point $|V_{e}^{\pm}|=|\mathcal{U}|$.
This critical threshold is central to our analogue model, marking the boundary between distinct topological phases that we identify with the horizon formation condition. The physical origin of this QPT can be understood by drawing a parallel to Bosonic systems with higher spatial derivatives, such as the $z=3$ model studied in Ref. Kumar and Shankaranarayanan [2017]. There, it was shown that higher-order derivative terms — corresponding to extended lattice neighbors — can violate the area law of entanglement entropy and trigger a phase transition due to the accumulation of zero modes. Similarly, in our model, the chirality operator acts as a higher-order correction to the standard $XX$ chain dynamics. While $H_{XX}$ governs nearest-neighbor transport, $H_{\chi}$ induces three-site correlations that effectively modify the dispersion relation’s topology. The critical point $|V_{e}^{\pm}|=|\mathcal{U}|$ can be mapped to a Lifshitz-like transition where the Fermi surface topology changes Ramazashvili [1999], Ardonne et al. [2004], Fradkin [2013], mimicking the causal disconnection of a horizon.
@@
This state encodes the thermal response of the qubit to the sudden formation of the chiral background. In the rest of this work, we analyze this as an analogue of HR in BH.
Analogue BH Geometry: To go about this, we take the continuum limit of the above spin-chain model. As shown in Ref. Jaiswal and Shankaranarayanan [2025], the spin chain maps onto a Dirac fermion propagating in a $(1+1)$-dimensional geometry:
@@
$ds^{2}=\,\Biggl\{\begin{array}[]{@{}ll@{}}dt^{2}-\frac{1}{\mathcal{U}^{2}(x)}dx^{2},&\,\,t<t_{0}\\ g(x)dt^{2}-2\sqrt{1-g(x)}\frac{dt\,dx}{\mathcal{U}(x)}-\frac{dx^{2}}{\mathcal{U}^{2}(x)},&\,\,t\geq t_{0}\end{array}~,$ (11)
where $g(x)=1-{V_{e}^{\pm}}^{2}/\mathcal{U}^{2}(x)$. Notably, the post-quench metric ($t\geq t_{0}$) corresponds to the Gullstrand-Painlevé coordinates of a spherically symmetric spacetime Shankaranarayanan [2003], confirming the emergence of a BH geometry.
Figure 2: Penrose diagram of the gravitational collapse process where a single shock wave located at $w=w_{0}$ collapse to form a BH.
To analyze BH, we employ null coordinates to track the mixing of positive and negative frequency modes. The “in” region is described by coordinates $(t,\mathcal{X})$, while the “out” region utilizes $(\tau,\mathrm{X})$; the explicit transformations are detailed in Table 1. For the spatial inhomogeneity, we adopt the profile Jaiswal and Shankaranarayanan [2025]:
where $\sigma$ controls the width of the mass distribution. Horizon formation occurs when the ”velocity” condition $|V_{e}^{\pm}|=|\mathcal{U}|$ is met. For widths below a critical value $\sigma_{c}\equiv 1/(2{V_{e}^{\pm}}^{2})$, this condition yields two symmetric horizons at:
Plane Waves
Gaussian Wave Packets
@@
where $\mathcal{X}_{h}=\pm\sigma\sqrt{2(\sigma_{c}-\sigma)}$. This logarithmic redshift mixes positive and negative frequency modes. Computing the Bogoliubov coefficients $\beta_{ff^{\prime}}$ via the Dirac inner product on a Cauchy surface yields the particle spectrum at late times:
$n_{f}^{P}=\int_{0}^{\infty}\!\!\!\!df^{\prime}|\beta_{ff^{\prime}}|^{2}=\frac{1}{e^{{f}/{T_{H}}}+1},\,T_{H}=\sqrt{\frac{\sigma_{c}-\sigma}{2\sigma^{2}\sigma_{c}^{2}\pi^{2}}}~.$ (15)
@@
where $\mathcal{D}_{-\nu}$ is the Parabolic cylinder function, $z=-2w_{h}-if^{\prime}$, and $\nu=1+if/(2\pi T_{H})$. Unlike the delocalized plane-wave case, this ratio is generally non-thermal. This non-thermal behavior can be understood in terms of grey-body factors describing the fact that not all modes produced near the horizon reach the detector unchanged: part of the wave packet is reflected, scattered, or distorted by the background before detection Unruh [1976a], Page [1976].
@@
Figure 3 illustrates the effects of localization at fixed $T_{H}=2$ as determined from (15) with $\sigma=0.1014$ and $\sigma_{c}=0.1115$ is fixed by choosing $v=1$, $g_{1}=1$, and $g_{2}=1/2$. The left panel shows the logarithmic Bogoliubov ratio. At low frequencies, it follows the linear thermal scaling, but at high frequencies, it saturates to a constant. This indicates that physical detectors are insensitive to trans-Planckian modes. The right panel confirms that the particle number spectrum $n_{f}^{G}$ deviates from the ideal thermal prediction Alkac and Ozgun [2025]. Thus, strict thermality is a feature of idealized infinite-duration measurements; realistic probes observe deviations at high energies.
@@
Finally, we examine the statistical nature of the radiation. Figure 4 displays the spectral distribution of outgoing modes for both probe types plotted against the Poisson parameter $s=\braket{n_{f}}$. Contrary to Wigner–Dyson statistics $P_{WD}=(\pi s/2)e^{-\pi s^{2}/4}$ Mehta [2004], both plane-wave and Gaussian distributions are well-described by the Poisson form $e^{-s/s_{0}}$. This demonstrates that the statistical character of HR is robust: even when localization effects modify the spectrum, the emission process remains fundamentally Poissonian.
@@
This is the key result of this work regarding which we want to discuss the underlying physical reason for the universal Poissonian behavior. HR originates from vacuum pair-production events which, in the absence of significant backreaction, are statistically independent and uncorrelated in time. This independence renders the emission a Poisson process, analogous to random tunneling events Shankaranarayanan et al. [2002], Shankaranarayanan [2003]. Since Gaussian wave packets are formed by linear superpositions of plane waves without introducing phase coherence, they preserve this intrinsic stochasticity. Thus, the Poissonian statistics reflect the fundamental randomness of vacuum fluctuations, which is robust against the spatial localization of the probe.
Qubit as a quantum probe: To operationalize the detection of HR, we couple a stationary qubit to the spin chain, acting as an Unruh-DeWitt detector Unruh [1976b]. The qubit’s relaxation and decoherence dynamics encode the correlation structure of the background field. The fidelity of this “thermometer” depends critically on the coupling strength, revealing two distinct physical regimes.

### v2

@@ Abstract
We investigate the emergence and detection of Hawking radiation (HR) in a 1D chiral spin chain model, where the gravitational collapse is simulated by a sudden quantum quench that triggers a horizon-inducing phase transition. While our previous work Jaiswal and Shankaranarayanan (2025) established that this model mimics black hole (BH) formation conditions even when the Hoop conjecture is seemingly violated, we here focus on the resulting stationary radiation spectrum and its detectability. By mapping the spin chain dynamics to a Dirac fermion in a curved $(1+1)$-dimensional spacetime, we analyze the radiation using two complementary approaches: field-theoretic modes and operational quantum sensors. First, using localized Gaussian wave packets to model realistic detectors, we find that the radiation spectrum exhibits deviations from the ideal Planckian form, while retaining robust Poissonian statistics that signal the loss of formation-scale information. Second, we introduce a qubit coupled globally to the chiral spin chain. Rather than acting as a localized point detector (which would break translational invariance), the qubit couples to the collective chiral excitations of the analogue spacetime, functioning as a global quantum sensor. We demonstrate that the decoherence dynamics of this macroscopic probe faithfully encode the thermal nature of the HR in the weak-coupling limit, while strong-coupling dynamics reveal non-Markovian memory effects driven by the horizon. These results provide a clear operational protocol for distinguishing genuine analog HR from environmental noise in quantum simulation platforms.
@@
Introduction: A central challenge in semiclassical and quantum gravity is the direct experimental verification of Hawking radiation (HR) Parker (1975); Hawking (1976); Corley and Jacobson (1996); Shankaranarayanan et al. (2002, 2001); Casadio (2002); Fedichev and Fischer (2003); Kerner and Mann (2006, 2008); Kumar and Shankaranarayanan (2016); Pal et al. (2022); Smerlak and Singh (2013); Pal and Fischer (2024). While astrophysical observations, such as strong-field gravitational lensing, provide critical probes of the near-horizon geometry of black holes (BHs) Mandal and Shankaranarayanan (2026); Virbhadra (2022), the quantum dynamics of the horizon itself remain obscured by the weakness and delocalization of the radiation. Analogue gravity platforms have emerged as a powerful tool, as they enable controlled access to near-horizon quantum effects in laboratory settings ranging from fluids to optical systems Fedichev and Fischer (2004); Barcelo et al. (2005); Unruh and Schutzhold (2012); Rajagopal et al. (2014); Steinhauer (2014); Steinhauer and de Nova (2017); Banerjee et al. (2019); Guo (2022).
@@
To extract these elusive signatures from the background of a many-body system, we turn to the paradigm of quantum sensing Degen et al. (2017); Pirandola et al. (2018); DeMille et al. (2024); Mercer et al. (2025); Zhang and Zhuang (2021). Controlled spin-1/2 particles (qubits) have established themselves as highly sensitive probes of complex quantum environments, underpinning modern sensing technologies across molecular, atomic, and solid-state systems Degen et al. (2017). When coupled to an external bath, a qubit’s decoherence dynamics and population relaxation are governed by the spectral properties of the environment Płodzień et al. (2025). This allows for the precise extraction of noise spectra Clerk et al. (2010); Young and Whaley (2012) and Hamiltonian parameters Degen et al. (2017) solely through measurements on the two-level probe.
@@
Motivated by these capabilities, we employ a qubit coupled globally to the chiral chain, to probe the analogous radiation in the chiral spin chain Cong et al. (2021). In our previous work, we established that the chiral spin chain Hamiltonian can mimic gravitational collapse, satisfying black hole formation conditions even when the Hoop conjecture is violated Jaiswal and Shankaranarayanan (2025). In the present work, we focus on the implications for HR. Specifically, we investigate how the qubit’s decoherence profile encodes the emergent thermal features of the horizon.
@@
Our analysis proceeds in two steps. First, we examine the radiation spectrum directly using both idealized plane waves and physically realistic Gaussian wave packets. We show that while plane waves recover the standard thermal spectrum, localized packets reveal deviations akin to greybody factors Birrell and Davies (1982); Wald (1995); Brout et al. (1995); Unruh (1976); Page (1976), challenging the notion of HR as information-free. Second, we determine the operational conditions for detecting this temperature. We demonstrate that a qubit acts as a faithful thermometer only in the weak-coupling regime, whereas strong coupling leads to thermalization with the bulk environment rather than the horizon. Finally, we discuss the statistical nature of the emission, showing that the radiation remains Poissonian regardless of the geometric “Hoop” history, suggesting a universal erasure of formation scale information in the Hawking regime.
@@
Model setup: To explore the dynamics of HR, we model the gravitational collapse as a sudden quantum quench Byju et al. (2023); Kaplanek and Burgess (2021); Bongs et al. (2023); Covey et al. (2025); Balatsky et al. (2025). We model the transition from a pre-horizon mass distribution to a BH using a time-dependent Hamiltonian $H(t)$. We assume the system initially evolves under a background Hamiltonian $H_{0}$ (representing the pre-collapse phase) and undergoes a sudden change at $t=t_{0}$ representing the formation of the horizon:
@@
where $\Theta(t-t_{0})$ is the Heaviside step function and $H_{h}$ denotes horizon-inducing dynamics and the probe interaction. In our analogue model, the pre-horizon regime is governed by the standard $XX$ spin chain, given by Jaiswal and Shankaranarayanan (2025):
$H_{0}\equiv H_{XX}=-\sum_{j=1}^{N}\frac{u(j,j+1)}{2}\left(\sigma_{j}^{x}\sigma_{j+1}^{x}+\sigma_{j}^{y}\sigma_{j+1}^{y}\right),$ (2)
where $\sigma_{j}^{\alpha}$ are Pauli operators at site $j$ satisfying periodic boundary conditions, and $u(j,j+1)\equiv u(|(j+1)-j|)$ is a site-dependent coupling constant depending only on the distance between nearest-neighbor sites. For $t\geq t_{0}$, the BH formation is captured by the activation of the chirality term $H_{\chi}$ and the coupling to an external probe $H_{I}$. The quench Hamiltonian is defined as $H_{h}=H_{\chi}+H_{I}$. The chirality term is:
$H_{\chi}=\sum_{j=1}^{N}(v/4)\,\chi_{j}~,\quad H_{I}=\sum_{j=1}^{N}(g_{1}\sigma_{q}^{z}+g_{2}\sigma_{q}^{x})\chi_{j}/4,$ (3)
where $\chi_{j}=\vec{\sigma}_{j}\cdot(\vec{\sigma}_{j+1}\times\vec{\sigma}_{j+2})$ is three-spin chirality operator Horner et al. (2023); Forbes et al. (2023) and $v$ is a dimensionless coupling constant. The chirality operator induces handedness by breaking left–right symmetry and produces an imbalance in spin excitations, which cannot be generated by the inhomogeneous $XX$ spin interactions alone. When activated by the quench, this term converts the inhomogeneity of the nearest-neighbor exchange coupling into an effective curvature in the emergent spacetime. To detect the thermal spectrum of HR, we couple the chain to a qubit via the interaction term $H_{I}$ Quan et al. (2006); Liu et al. (2009); Jaiswal et al. (2022). Here $\sigma_{q}^{x,z}$ act on the qubit and $g_{1,2}$ are coupling strengths. It is crucial to note the physical nature of the interaction Hamiltonian $H_{I}$. The qubit operators are summed over all lattice sites $N$, meaning the qubit interacts with the total macroscopic chirality of the chain, $\sum_{j}\chi_{j}$. Physically, this implies that the qubit does not act as a localized defect or a point-like Unruh-DeWitt detector, which would inevitably break the translational invariance of the lattice. Instead, the qubit is uniformly coupled to the bulk, acting as a global quantum sensor that probes the collective zero-momentum field excitations of the emergent curved spacetime. This global coupling preserves momentum conservation, allowing for an exact analytical diagonalization while capturing the macroscopic particle production generated by the horizon. As shown in Fig. (1), coupling allows the qubit to exchange energy with the chain, inducing decoherence that serves as a proxy for the thermal bath of HR.
Figure 1: Schematic diagram of the qubit–environment coupling for an $N=8$ chiral spin chain: the qubit (red) couples to the spins (blue) via the chirality operator $\chi_{j}$, while the spins interact through XXXX and chiral terms.
@@
In each sector corresponding to eigenvalues $\pm\sqrt{g_{1}^{2}+g_{2}^{2}}$, the spin chain evolves under an effective Hamiltonian. Following the standard Jordan-Wigner and Fourier transformations Jaiswal and Shankaranarayanan (2025); Jaiswal et al. (2021), the total Hamiltonian is diagonalized in momentum space (where $p\in[-\pi,\pi)$):
@@
$E_{\pm}(p)=-2\,\mathcal{U}\cos(p)+V_{\cal E}^{\pm}\sin(2p)~,$ (6)
where $\mathcal{U}e^{ip}$ is the Fourier transform of $u(|(j+1)-j|)=\mathcal{U}$ (see Appendix A for details) and the effective chiral coupling is modified by the probe interaction: $V_{\cal E}^{\pm}=v\pm\sqrt{g_{1}^{2}+g_{2}^{2}}$.
Quantum Phase Transition and Time Evolution: Analysis of the ground-state energy density reveals a Quantum Phase Transition (QPT) driven by the competition between the nearest-neighbor hopping and the chiral interaction. A discontinuity in the second derivative of the energy density with respect to the effective coupling parameter $V_{\cal E}^{\pm}$ occurs at $|\mathcal{U}|=|V_{\cal E}^{\pm}|$, signaling a continuous QPT that separates the chiral and non-chiral phases.
This critical threshold is central to our model, marking the boundary between distinct topological phases that we identify with the horizon formation condition. The physical origin of this QPT can be understood by drawing a parallel to Bosonic systems with higher spatial derivatives, such as the $z=3$ model studied in Ref. Kumar and Shankaranarayanan (2017). There, it was shown that higher-order derivative terms — corresponding to extended lattice neighbors — can violate the area law of entanglement entropy and trigger a phase transition due to the accumulation of zero modes. Similarly, in our model, the chirality operator acts as a higher-order correction to the standard $XX$ chain dynamics. While $H_{XX}$ governs nearest-neighbor transport, $H_{\chi}$ induces three-site correlations that effectively modify the dispersion relation’s topology. The critical point $|V_{\cal E}^{\pm}|=|\mathcal{U}|$ can be mapped to Lifshitz-like transition where the Fermi surface topology changes Ramazashvili (1999); Ardonne et al. (2004); Fradkin (2013), mimicking the causal disconnection of a horizon.
@@
$|\Psi(t)\rangle=c_{+}|I_{+}\rangle\otimes\left(\prod_{p}|\psi_{E_{+,p}}(t)\rangle\right)+c_{-}|I_{-}\rangle\otimes\left(\prod_{p}|\psi_{E_{-,p}}(t)\rangle\right),$ (8)
@@
Analogue BH Geometry: To understand the emergence of the analogue event horizon, we must map the spin chain dynamics to a geometric continuum limit. A sudden quantum quench injects energy into the system, as the pre-quench ground state is no longer an eigenstate of the post-quench Hamiltonian. This non-adiabatic transition inevitably generates a superposition of excited states, manifesting physically as a directed flow of quasi-particles (a flux of excitations) that propagate ballistically outward Calabrese and Cardy (2016); Byju et al. (2023). In the geometric language of analogue gravity, this sudden quench is mathematically equivalent to the gravitational collapse of a null shell of matter. The propagating quasi-particles correspond to the null energy flux that separates the spacetime into two distinct regions governed by different effective metrics, dynamically inducing an effective event horizon Abajo-Arrastia et al. (2010). The HR we derive arises from the propagation of vacuum fluctuations with respect to this new causal structure, independent of the background excitations flowing into the black hole. As shown in Ref. Jaiswal and Shankaranarayanan (2025), the spin chain maps onto a Dirac fermion propagating in a $(1+1)$-dimensional geometry:
@@
$ds^{2}=\,\Biggl\{\begin{array}[]{@{}ll@{}}dt^{2}-d\mathcal{X}^{2},&\,\,t<t_{0}\\ g(x)dt^{2}-2\sqrt{1-g(x)}dt\,d\mathcal{X}-d\mathcal{X}^{2},&\,\,t\geq t_{0}\end{array}~,$ (10)
where $d\mathcal{X}=dx/\mathcal{U}(x)$ and $g(x)=1-{V_{\cal E}^{\pm}}^{2}/\mathcal{U}^{2}(x)$. Although, the nearest-neighbor exchange coupling $\mathcal{U}$ is an inhomogeneous quantity in both the pre and post-quench states, in the pre-quench regime this coupling merely rescales the local propagation velocity and can be absorbed into a coordinate redefinition, leaving the effective spacetime flat in $(t,\mathcal{X})$ coordinates. The quench, however, activates the chiral interaction, which explicitly breaks left–right symmetry and generates an imbalance in spin excitations, closely analogous to the charge separation underlying the chiral magnetic effect. This chiral term couples nontrivially to the existing inhomogeneities, and as a consequence the post-quench dynamics is no longer equivalent to that of flat space: the combined action of chirality and inhomogeneous $XX$ couplings curves the emergent spacetime and drives the system toward a topological QPT. The post-quench metric ($t\geq t_{0}$) corresponds to the Gullstrand-Painlevé coordinates of a spherically symmetric spacetime Shankaranarayanan (2003), with the spatial profile $\mathcal{U}$ acts as a position-dependent ‘speed of light’ for the emergent Dirac fermions in the engineered BH geometry.
Figure 2: Penrose diagram of the collapse process where a shock wave located at $w=w_{0}$ collapses to form a BH.
@@
To analyze BH, we employ null coordinates to track the mixing of positive and negative frequency modes. The “in” region is described by coordinates $(t,\mathcal{X})$, while the “out” region utilizes $(\tau,\mathrm{X})$; the explicit transformations are detailed in Table 1. For the spatial inhomogeneity, we adopt the profile Jaiswal and Shankaranarayanan (2025):
where $\sigma$ controls the width of the mass distribution. Horizon formation corresponds to $|V_{\cal E}^{\pm}|=|\mathcal{U}|$. For widths below a critical value $\sigma_{c}\equiv 1/(2{V_{\cal E}^{\pm}}^{2})$, this condition yields two symmetric horizons at:
Plane Waves
Gaussian Wave Packets
@@
where $\mathcal{X}_{h}=\pm\sigma\sqrt{2(\sigma_{c}-\sigma)}$. This logarithmic redshift mixes positive and negative frequency modes. The radiation spectrum is governed by the Bogoliubov coefficients $\alpha$ and $\beta$, which quantify the nontrivial mixing between positive and negative-frequency modes due to the time-dependent background (see Appendix B for their rigorous definitions via the Dirac inner product). Computing the Bogoliubov coefficients yield the particle spectrum at late times:
$n_{f}^{P}=\int_{0}^{\infty}\!\!\!\!df^{\prime}|\beta_{ff^{\prime}}|^{2}=\frac{1}{e^{{f}/{T_{H}}}+1},\quad T_{H}=\sqrt{\frac{\sigma_{c}-\sigma}{2\sigma^{2}\sigma_{c}^{2}\pi^{2}}}~.$ (14)
@@
where $\mathcal{D}_{-\nu}$ is the Parabolic cylinder function, $z=-2w_{h}-if^{\prime}$, and $\nu=1+if/(2\pi T_{H})$. Unlike the delocalized plane-wave case, this ratio is generally non-thermal.
@@
Figure 3 illustrates the effects of localization at fixed $T_{H}=2$ as determined from (14) with $\sigma=0.1014$ and $\sigma_{c}=0.1115$ is fixed by choosing $v=1$, $g_{1}=1$, and $g_{2}=1/2$. The left panel shows the logarithmic Bogoliubov ratio. At low frequencies, it follows the linear thermal scaling, but at high frequencies, it saturates to a constant. This indicates that physical detectors are insensitive to trans-Planckian modes. The right panel confirms that the particle number spectrum $n_{f}^{G}$ deviates from the ideal thermal prediction Alkac and Ozgun (2025). Thus, strict thermality is a feature of idealized infinite-duration measurements; realistic probes observe deviations at high energies.
@@
Finally, we examine the statistical nature of the radiation. To systematically analyze the statistical properties of the emitted radiation and construct the distributions shown in Fig. 4, we examine the level spacing statistics of the frequency modes. We first define a normalized probability density function for both the Plane ($P$) and Gaussian ($G$) wavepackets as:
$\mathcal{P}_{P,G}(f)={n_{f}^{P,G}}\Biggm/{\int_{0}^{\infty}df\,n_{f}^{P,G}}~,$ (16)
where $n_{f}^{P,G}$ represents the respective frequency spectrum (defined in Eq. (14) for the plane-wave case). Using this probability density, a sequence of frequency samples is numerically generated and sorted in ascending order. From these ordered samples, we compute the consecutive frequency spacing variable, $s\equiv\langle n_{f}^{P,G}\rangle$. The quantity $\mathcal{N}^{P,G}$ plotted on the vertical axis of Fig. (4) is the numerically estimated probability density of this spacing variable $s$. It is constructed by histogramming the computed $s$ values and normalizing the area of the histogram to unity. Consequently, $\mathcal{N}^{P,G}$ does not possess a predefined analytic expression but is purely derived from the simulated spectral data.
To interpret these numerical results, we compare $\mathcal{N}^{P,G}$ against two standard analytical references. As clearly seen in Fig. (4), the numerical data rigorously follows the analytic Poissonian distribution, $P(s)=e^{-s/s_{0}}$ (solid green line), which is the expected behavior for uncorrelated thermal emission. It distinctly deviates from the Wigner-Dyson distribution Mehta (2004), $P_{WD}(s)=(\pi s/2)e^{-\pi s^{2}/4}$ (dashed magenta line), which would otherwise indicate correlated level repulsion.
@@
Contrary to Wigner–Dyson statistics, both plane-wave and Gaussian distributions are well-described by the Poisson form $e^{-s/s_{0}}$. This demonstrates that the statistical character of HR is robust: even when localization effects modify the spectrum, the emission process remains fundamentally Poissonian.
@@
This is the key result of this work regarding which we want to discuss the underlying physical reason for the universal Poissonian behavior. HR originates from vacuum pair-production events which, in the absence of significant backreaction, are statistically independent and uncorrelated in time. This independence renders the emission a Poisson process, analogous to random tunneling events Shankaranarayanan et al. (2002); Shankaranarayanan (2003). Since Gaussian wave packets are formed by linear superpositions of plane waves without introducing phase coherence, they preserve this intrinsic stochasticity. Thus, the Poissonian statistics reflect the fundamental randomness of vacuum fluctuations, which is robust against the spatial localization of the probe.
From the perspective of the microscopic spin chain, the distinction between these modes is physically interesting. Plane waves correspond to highly idealized, globally delocalized momentum eigenstates of the lattice. In contrast, Gaussian wavepackets represent physically realizable, localized spin excitations. While the kinematic existence of the horizon is fixed by the effective metric, our results demonstrate that the exact thermality of HR is an artifact of idealized plane waves. Localized wavepackets which represent the actual excitations a finite detector would absorb–inherently retain non-thermal correlations.
It is important to note that the fundamental scattering dynamics of plane waves and the resulting greybody factors are well-established results in the standard literature of quantum field theory in curved spacetime Birrell and Davies (1982); Wald (1995); Brout et al. (1995); Unruh (1976); Page (1976). Our motivation for explicitly comparing plane-wave and Gaussian-wavepacket descriptions is not to present novel greybody physics, but rather to address the specific operational requirements of an analog quantum simulator. In this framework, plane-wave solutions are utilized to maintain analytic control and to unambiguously extract the underlying thermal Hawking spectrum. In contrast, the Gaussian-wavepacket formalism is necessary to faithfully model the finite-time, spatially localized excitation and detection protocols that are inherent to a realistic spin-chain experiment. This dual approach bridges the idealized analytical thermality of the BH horizon with the practical realities of localized measurements in condensed matter systems.
Qubit as a quantum probe: Because the qubit couples to the global modes of the spin chain, its response reflects the macroscopic entanglement and particle production generated by the dynamical quench. In the weak-coupling (Markovian) limit, the qubit acts as a collective thermometer, relaxing to a steady state that faithfully registers the Hawking temperature of the bulk radiation. In contrast, the strong-coupling regime is dominated by measurement back-action and non-Markovian revivals, masking the purely thermal signature but revealing the robust entanglement structure of the post-quench vacuum.